# Observability of dynamical tides in merging eccentric neutron star binaries

### v1

@@ III.5 Detectability of the tidal oscillations
However we caution that the assumed detectability limit of SNR$=8$ is usually assumed for GW sources with available matched filtering templates, but it is unclear whether a matched filtering search is feasible for NS binaries with dynamical tides. The detection and measurement of NSs with dynamical tides may be more difficult.
@@ III.6 Implications for the equation of state
While dynamical tides affect the evolution of circular binaries as well, they only become significant very close to the merger Hinderer et al. (2016); Schmidt and Hinderer (2019). Therefore, the tidal phase shift is mostly determined by adiabatic tides, which only depend on the mass-weighted tidal deformability:
@@
Hence, the measurement of adiabatic tides only provide information about this specific combination of binary parameters, while $\Lambda_{1}$ and $\Lambda_{2}$ could only be constrained by imposing additional constraints on the EoS. On the other hand, dynamical tides also carry information about the f-mode frequency of tidal oscillations. This could resolve the degeneracy between the binary parameters and could be used together with quasi-universal relations to determine $\Lambda_{1}$ and $\Lambda_{2}$ independently, thus also providing information about the radii, $R_{1}$ and $R_{2}$ of the two NSs.
@@ Appendix B Equations of motion
$e=\sqrt{1+\frac{2E_{\mathrm{orb}}L_{\mathrm{orb}}^{2}}{m_{\mathrm{tot}}^{2}}}\>$ $,\quad a=\frac{m_{\mathrm{tot}}}{2E_{\mathrm{orb}}}\>,$ (59)

### v2

@@ III.4 Comparison to the literature
Another interesting phenomenon was shown in Ref. Yu et al. (2024), who found that initially circular binaries can become slightly eccentric due to strong tidal torques during the late inspiral. They showed that for binaries of highly spinning NSs with anti-aligned spins the eccentricity can become as large as $0.05-0.1$ right before the merger. Fig. 9 shows results from our simulations on the eccentricity evolution of the binary without the 1PN and 2PN terms. Similarly to Ref. Yu et al. (2024) we show the evolution of the osculating orbital element $ec_{\varphi}$ compared to the point-particle (pp) case, where
$ec_{\varphi}\equiv e\cos(\varphi-\varphi_{0})=\frac{p}{r}-1=\frac{r^{3}\dot{\varphi}^{2}}{m_{\mathrm{tot}}}-1\>,$ (35)
with $\varphi_{0}$ being the argument of pericenter, and $p$ the semilatus rectum. Our findings are consistent with those of Ref. Yu et al. (2024) in that a finite eccentricity is induced by tides right before the merger. We find that this effect is substantial even for non-spinning NSs. However, we note that eccentricities of $e\gtrsim 0.01$ are only induced $2-4$ orbits before direct collision of the NSs, where the Newtonian approximation might break down.
Figure 9: The change in the oscillatory eccentricity $ec_{\varphi}$ due to tidal interactions compared to the point-particle case. The solid blue and yellow lines correspond to non-spinning NSs with different tidal deformabilities, while the red line correspond to NSs with the f-mode frequency modified to $900$ Hz, representing highly-spinning NSs (with $f_{\mathrm{rot}}\approx 800$ Hz), with anti-aligned spins. Each vertical dashed line represents a single orbit.
@@ III.5 Detectability of the tidal oscillations
However we note that the assumed detectability limit of SNR$=8$ is usually assumed for GW sources with available matched filtering templates, but it is unclear whether a matched filtering search is feasible for NS binaries with dynamical tides. The detection and measurement of NSs with dynamical tides may be more difficult.
@@ III.6 Implications for the equation of state
While dynamical tides affect the evolution of circular binaries as well, they only become significant very close to the merger Hinderer et al. (2016); Schmidt and Hinderer (2019). Therefore, the tidal phase shift is mostly determined by adiabatic tides, which, to the leading 5PN order, only depend on the mass-weighted tidal deformability:
@@
Hence, the measurement of adiabatic tides only provide information about this specific combination of binary parameters to the leading 5PN order, while $\Lambda_{1}$ and $\Lambda_{2}$ could only be constrained by imposing additional constraints on the EoS. On the other hand, dynamical tides also carry information about the f-mode frequency of tidal oscillations. This could resolve the degeneracy between the binary parameters and could be used together with quasi-universal relations to determine $\Lambda_{1}$ and $\Lambda_{2}$ independently, thus also providing information about the radii, $R_{1}$ and $R_{2}$ of the two NSs.
@@
It needs to be added here that on top of the 5PN leading-order contributions, higher-order terms also contribute to adiabatic tides. The next-to-leading-order contribution to the GW phase shift comes from the 6PN effects, which scale with $\delta\psi\propto\delta\Lambda\omega_{\mathrm{orb}}^{7/3}$ Vines et al. (2011); Vines and Flanagan (2013); Jiménez Forteza et al. (2018), where
$\delta\Lambda$ $=\left(\frac{5095}{28}-\frac{15895}{28\chi_{1}}+\frac{5715\chi_{1}}{14}-\frac{325\chi_{1}^{2}}{7}\right)\chi_{1}^{5}\Lambda_{1}$
+(12)+(1$\leftrightarrow$ 2)\>,+ ( 1 ↔ 2 ) , (46)
and $\chi_{A}=m_{A}/m_{\mathrm{tot}}$. While the phase shift induced by the 6PN term is approximately an order of magnitude lower than that of the 5PN term, it still reaches values of $\mathcal{O}(0.1-1)$ during the late inspiral, making it comparable to the dynamical tides considered in this work close to merger. Furthermore, since the phase shift is proportional to $\delta\Lambda$, which is linearly independent of $\tilde{\Lambda}$ as a function of $m_{1}$ and $m_{2}$, one might also hope to break the degeneracy between $\Lambda_{1}$ and $\Lambda_{2}$ by measuring the 6PN effect. However, note that for a binary NS with $m_{1}=m_{2}=m$, the 5PN and 6PN tidal coefficients (Eqs. 45 and III.6) simply become $\tilde{\Lambda}=\Lambda_{1}+\Lambda_{2}$ and $\delta\Lambda\propto(\Lambda_{1}+\Lambda_{2})=\tilde{\Lambda}$, respectively. Consequently, for equal-mass binaries, the adiabatic tidal contributions remain degenerate at least up to the 6PN order for binaries with the same $\tilde{\Lambda}$ value. In the non-equal mass case shown in Fig. 13 for the same values of tidal deformabilities, we find that the largest relative difference in the 6PN phase shift is $\sim 1\%$, corresponding to a phase difference of $\mathcal{O}(10^{-3}-10^{-2})$. Therefore, we conclude that 6PN contributions are typically insufficient to break the degeneracy in the measurement of tidal deformabilities with the current sensitivity of GW detectors.
@@ Appendix B Equations of motion
$e=\sqrt{1+\frac{2E_{\mathrm{orb}}L_{\mathrm{orb}}^{2}}{m_{\mathrm{tot}}^{2}}}\>$ $,\quad a=-\frac{m_{\mathrm{tot}}}{2E_{\mathrm{orb}}}\>,$ (61)